OpenStudy (anonymous):

The area of a regular hexagon in 65 in (squared). find the length of a side. Round to the nearest tenth. a. 6.6 in. b. 5 in. c. 25 in. d. 8.7 in .

OpenStudy (anonymous):

Area = (3√3 s2)/ 2 (3√3 s2)/ 2 =65 s=5.01 nearly 5 in

OpenStudy (anonymous):

How did you do that?

OpenStudy (anonymous):

just calculate it 3√3 s2 =130 s2=130/3√3 s2=25.019 s=5.001

OpenStudy (anonymous):

Thanks